OpenStudy (anonymous):

what is this question asking for: Determine all real values t and s (1-s,3)=(7,t2)?

OpenStudy (anonymous):

Well, on the face of it, s = -6 and t= sqrt(3).

OpenStudy (anonymous):

\[(1-s,3)=(7,t^2)\] \[1-s=7\]\[3=t^2\] \[s=-6\]\[t=\sqrt{3}\]

OpenStudy (anonymous):

wait, sorry, it's t-2. *facepalm*

OpenStudy (anonymous):

In this case \[t-2=3\]\[t=5\]

OpenStudy (anonymous):

Thank you so very much, it is greatly appreciated